Problem Given Vo in the network, find I 1 kn 1 kn 1 kn 12 V 2 kn Vo 4V 6V 1 km ww
Answer #1

Given that Given Vo in the netook, incd IA Aasa t イゴV Vo 4V Adying KvL V」= 2+4 ニ16V。 loop-2, Q= n+nt 9- V2ニ6-V」 =6-16 = -loニー22へ 16 22x 16 へてて- * Vu こ12-22 -10 20V Iy Vy -20mA / 1に kcL -20x 1o 2 2 X 1o I A A 2 mA CS Scanned with CamScanner